The electric potential is described by a function V=10xy-6x^2+4y^3z. Calculate the electric field at the point [3.00,-2.00,5.00], write it in vector form.
Answer #1

V at [3.00,-2.00,5.00] = 10*3*-2 - 6*9 + 4*-8*5 = -274 V

didtance d = sqrt(3^2 + 2^2 +5^2) = 6.16

E = V/r = -274/6.16 = -44.44 N/C